# Exact mobility edges in quasiperiodic network models with slowly varying potentials

### v1

@@ V Conclusions
We propose a new class of 1D quasiperiodic network models with exact MEs, which are featured by slowly varying potentials. Unlike the previously proposed quasiperiodic network models [35], this new class doesn’t exhibit hidden self-duality, since the potential $\cos(\pi\alpha i^{\nu})$ is not self-dual. To obtain the MEs of quasiperiodic network models with slowly varying potentials, we present two distinct methods. The first involves integrating out the periodic sites to obtain an effective model, where $g(E)V$ represents the effective potential and $f(E)$ denotes the effective eigenenergy. Substituting $f$ and $g$ into $f=\pm(2t\pm gV)$ gives the concrete MEs. It should be noted that $f=\pm(2t\pm gV)$ is not guaranteed by the hidden self-duality, but comes from the MEs of the original slowly varying models. Additionally, we present a more convinced method by connecting the real-complex transition points of the eigenvalue equations and the localized-delocalized transition points of the quasiperiodic chain, which yields results in excellent agreement with the results of the first method.
Now, it is quite possible that we are in a position to reflect on the possible future of quantum simulations. In this work, we show that the rich phases in the quasiperiodic network models can be simulated using some simulation platforms, including optical and acoustic waveguides [63, 64, 65], ultracold atoms, superconducting qubits [66, 67, 68, 69, 70, 71], in which most of the important physics can be obtained using some small quantum systems. In recent years, these small systems have been used to demonstrate these fundamental physics. Yet, the detailed physics, including the scaling laws and the phase boundaries has not been determined. This is different from that in the community of condensed matter physics, in which the phase transition and the associated scaling exponents can be precisely determined in experiments. The similar accuracy has not yet been achieved in quantum simulation. In the future, with the increasing of experimental techniques, it is quite possible that these simulating platforms can be used to determine the precise phase boundary and the associated critical exponents, making quantum simulations a powerful tool to explore possible new physics. At the present stage, only some intriguing phenomena have been simulated. If this kind of accuracy has been achieved, the physical models may serve as an important platform for exploring Anderson transition in the network models [72], which will help us to resolve the long-standing problems in this physics. Especially, in the presence of many-body interaction [73, 74, 75] and in the higher dimensional models with various potentials [76, 77], the physics cannot be efficiently simulated using classical computers, and these possible new physics may be simulated using these intriguing platforms.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362) and the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500). A.-M. G. is supported by the NSFC (Grant Nos. 12274466 and 11874428) and the Hunan Provincial Science Fund for Distinguished Young Scholars (Grant No. 2023JJ10058).

### v2

@@ V Conclusions
To conclude, in this work we propose a new class of 1D quasiperiodic network models slowly varying potentials and we can exactly determine their MEs. Unlike the previously proposed quasiperiodic network models [28], this model doesn’t exhibit hidden self-duality, since the slowly varying potential is not self-dual. We integrate out the periodic sites to obtain an effective model, where $g(E)V$ represents the effective potential and $f(E)$ denotes the effective eigenenergy. Substituting $f$ and $g$ into $f=\pm(2t^{\kappa}\pm gV)$ will give the concrete MEs, following Das Sarma et. al. [59]. We confirm these results using numerical method. With this method, we can study the exact MEs with slowly varying potentials in some much more complicated models, including non-Hermitian ones. In any way, the exact MEs in models with slowly varying potentials should greatly broaden our understanding of MEs in 1D disordered models.
Now, it is quite possible that we are in a position to reflect on the possible future of quantum simulations. In this work, we show that the rich phases in the quasiperiodic network models can be simulated using the optical waveguides. In actually, there are much more platforms to simulate these models, including optical and acoustic waveguides [75, 33, 76], ultracold atoms, superconducting qubits [77, 78, 79, 80, 81, 82], in which most of the essential physics in these models can be obtained using some small quantum systems. In recent years, these small systems have been used to demonstrate these fundamental physics. Yet, the detailed physics, including the scaling laws and the phase boundaries have not been precisely determined. This is different from that in the community of condensed matter physics, in which the phase transition and the associated scaling exponents have been precisely determined in experiments. The similar accuracy has not yet been achieved in quantum simulation. In the future, with the increasing of experimental techniques, it is quite possible that these simulating platforms can be used to determine the precise phase boundary and the associated critical exponents, making quantum simulations a powerful tool to explore possible new physics. At the current stage, only some intriguing phenomena have been simulated. If this kind of accuracy can be achieved, the physical models may serve as an important platform for exploring Anderson transition in the network models [83], which will help us to resolve the long-standing problems in this physics. Especially, in the presence of many-body interaction [84, 85, 86] and in the higher dimensional models with various potentials [87, 88], these physics cannot be efficiently simulated using classical computers, however, it is quite possible that these new physics can be simulated using these intriguing platforms.
@@ Acknowledgements.
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500000), the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362) and the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500). A.-M. G. is supported by the NSFC (Grant Nos. 12274466 and 11874428) and the Hunan Provincial Science Fund for Distinguished Young Scholars (Grant No. 2023JJ10058).